OpenStudy (anonymous):

f(x)=(x-4)(2x-1)(x+7)

OpenStudy (anonymous):

f(x)=(x-4)(2x-1)(x+7) = (x-4)( 2x^2 + 13x - 7) = ( 2x^3 + 13x^2 - 7x) - ( 8x^2 + 52x - 28) = 2x^3 + 5x^2 - 59x + 28